### v1

# Sequential Change Detection under Markov Setup with Unknown Prechange and Postchange Distributions

Abstract
In this work we extend the results developed in 2022 for a sequential change detection algorithm making use of Page’s CUSUM statistic,empirical distribution as an estimate of pre-change distribution and a universal code as a tool for estimating the post change distribution, from IID to Markov setup.
I Introduction
Change detection has numerous applications including but not limited to industrial quality control,fault detection in computer communication networks and regime shift detection in financial markets.
Wald formalized the sequential hypothesis testing problem in 1945 [1].He proposed the Sequential Probability Ratio Test(SPRT) and proved its optimality along with Wolfowitz[2].Lai proved asymptotic optimality of invariant SPRT for classes of sources that satisfy a stability condition for likelihood ratio [12].
Page introduced the Cumulative Sum Test (CUSUM) to address the problem of sequential change detection in 1954 [4].Lorden proved the asymptotic optimality of CUSUM for memoryless sources in minimax framework [6] while Moustadikes proved its exact optimality [15]. Shiryaev [5, 10] introduced a different approach to the change detection problem under Bayesian framework which has been extensively studied in literature.Lorden’s result was extended for sources with memory by Bansal and Kazakos [16] who assume independence of observations before and after change which was further weakened in Lai’s extension [20].[8, 21] are other extensions of CUSUM beyond i.i.d. framework.
Page’s CUSUM Test requires that both pre-change and post-change distributions should be known which is not always the case in real world situations.Jacob and Bansal [24] used ideas from information theory to develop a modified SPRT test as well as modified CUSUM test for the case where post-change distribution is unknown and a universal code is used in the place of log-likelihood of post-change distribution for memoryless sources.They showed the optimality of the modified test in asymptotic regime where the rate of false alarm goes to 0.They extended the idea of modified test for sources with memory under Berk’s mixing conditions [25].
Malik and Bansal [26] considered the case where both pre-change and post-change distributions are unknown for memoryless sources.They further modified the test introduced by Jacob and Bansal [24] ,by using an empirical estimate in the place of true prechange distribution without affecting the post-change distribution estimate and proved the asymptotic optimality of the proposed test.
In this paper, we extend the modification demonstrated by Malik and Bansal [26] for the more general class of stationary and ergodic finite-order Markov sources that satisfy Berk’s mixing conditions[7] and make use of Bansal and Kazakos’ extension [16] to study the performance of proposed test and establish its optimality in the asymptotic regime where the rate of false alarm goes to 0.
Our main contributions in the paper are:
1.Upper bound on probability of error for proposed test in Lemma 3.
2.Showed that the proposed test will terminate definitely under alternate hypothesis in Lemma 4.
3.Evaluated the performance of proposed test by studying linear cost function of the number of samples required to achieve a specific probability of error in Theorem 2 and the linear cost function of detection delay required to achieve a specific false alarm rate in Theorem 5.
4.Established the optimality of proposed test in the asymptotic regime in Theorem 6.
Section II introduces the preliminaries required for the rest of the paper.In Section III, we introduce the empirical estimate,the modified test and study behaviour of proposed test under pre-change and post-change distribution.In Section IV, we study the performance of modified test and prove its asymptotic optimality.
II Preliminaries
Let $X=\{X_{n}\}_{n=1}^{\infty}$ be a source with finite alphabet $\mathcal{X}$.A sequence of source symbols $x_{i},x_{i+1},\dots,x_{j}$ is denoted by $x_{i}^{j}$ and corresponding random variables $X_{i},X_{i+1},\dots,X_{j}$ is denoted by $X_{i}^{j}$.All the logarithms will be taken to the base 2 in the remaining, unless specified otherwise.
A fixed-to-variable code (FV code) is a map $\phi_{n}:X^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length strings.Let $L(x_{1}^{n})=|\phi_{n}(x_{1}^{n})|$ denotes the length of the codeword for a given sequence $x_{1}^{n}$ which satisfies prefix condition.Therefore, Kraft inequality holds, given as
$\sum_{x_{1}^{n}\in X^{n}}2^{-L(x_{1}^{n})}\leq 1$
Let $\mu$ be the distribution of $X=\{X_{n}\}_{n=1}^{\infty}$.The entropy rate for stationary $\mu$ is
$H(\mu)=-\lim_{n\to\infty}\frac{1}{n}\sum_{x_{1}^{n}\in X^{n}}\mu(x_{1}^{n})\log\mu(x_{1}^{n})$
For a stationary ergodic source, Shannon-Mcmillian-Brieman theorem states
$\lim_{n\to\infty}\frac{1}{n}\log\mu(x_{1}^{n})=H(\mu)\quad\mu\text{ a.s.}$
A FV code with length function $L(x_{1}^{n})$ is said to be pointwise universal if
$\frac{1}{n}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\to 0\quad\mu\text{ a.s.}$
For stationary ergodic sources, the existence of such codes is guaranteed [22, 23].For a class of sources $\mathcal{M}$, a FV code with length function $L(x_{1}^{n})$ is said to be strongly pointwise universal if:
$R_{L}(n)=\sup_{\mu\in\mathcal{M}}\max_{x_{1}^{n}\in X^{n}}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\sim o(n)$
The condition for pointwise universality (2.4) is equivalent to having $\mu(J_{\epsilon}<\infty)=1$ for all $\epsilon>0$, where
$J_{\epsilon}=\sup_{n\geq 1}\left\{n:\left|\frac{1}{n}\left(L(x_{1}^{n})+\log\mu(x_{1}^{n})\right)\right|\geq\epsilon\right\}$ (2.6)
A stronger version of the convergence condition (2.4) requires that
$\mathbb{E}_{\mu}(J_{\epsilon})<\infty\quad\text{for all}\quad\epsilon>0$ (2.7)
Strongly pointwise universal (2.5) codes satisfy this condition which is called 1-quick universal by Jacob [25], for every source in the class $\mathcal{M}$.For Markov sources upto a fixed order, the existence of codes that satisfy all these conditions are known(Lempel-Ziv codes (LZ-77, LZ-78) [9, 11] and the Context Tree Weighting code) [19] and we consider such codes only in our paper.
Theorem 1(Barron)[13, 18]: Let $\mu$ be a stationary ergodic process and $\nu$ be a stationary Markov process of finite order such that $\mu\ll\nu$.Then
$\frac{1}{n}\log\frac{\mu(x_{1}^{n})}{\nu(x_{1}^{n})}\to D(\mu||\nu)\quad\mu\text{ a.s.}\quad$
where $D(\mu||\nu)$ is the divergence rate.
The pointwise convergence in the above theorem is equivalent to having $\mu(K_{\epsilon}<\infty)=1\quad\text{for all}\quad\epsilon>0$ where
$K_{\epsilon}=\sup_{n\geq 1}\left\{n:\left|\frac{1}{n}\log\frac{\mu(x_{1}^{n})}{\nu(x_{1}^{n})}-D(\mu||\nu)\right|\geq\epsilon\right\}\quad$
A stronger version of this convergence, called 1-quick convergence of log-likelihood ratios according to Lai’s terminology [12], requires that
$\mathbb{E}_{\mu}(K_{\epsilon})<\infty\quad\text{for all}\quad\epsilon>0.\quad$
For a fixed sample hypothesis test with $H_{0}:\mu=\mu_{0}$ as the null hypothesis and $H_{1}:\mu=\mu_{1}$ as the alternative, consider the discriminant function
$h(x_{1}^{n})=\log\mu_{1}(x_{1}^{n})-\log\mu_{0}(x_{1}^{n})-n\lambda$
After observing $n$ samples $x_{1}^{n}$, we reject the null hypothesis if $h(x_{1}^{n})\geq 0$. The following lemma characterizes the performance of this test.
Lemma 1 [3, 17]:If $\mu_{0}$ and $\mu_{1}$ are stationary Markov process of finite order and stationary ergodic process such that $\mu_{1}\ll\mu_{0}$ and $D(\mu_{1}\|\mu_{0})>\lambda$. For the test induced by the discriminant function $h(x_{1}^{n})$ we have:
$\mu_{0}\big(h(x_{1}^{n})\geq 0\big)$  $\;\leq\;2^{-n\lambda}$ (2.12)
$\lim_{n\to\infty}\mu_{1}\big(h(x_{1}^{n})\geq 0\big)$ $=1.$ (2.13)
Equation (2.13) remains true even if the threshold 0 is replaced by any other constant.The lemma follows immediately from the Shannon–McMillan–Breiman theorem.
Berk’s Mixing Conditions [7, 16, 25]: Let
$p_{n}(\Delta)=\mu_{1}\left(\frac{1}{n}\log\frac{\mu_{1}(x_{1}^{n})}{\mu_{0}(x_{1}^{n})}<\Delta\right)$
and
$q_{n}(\Delta)=\mu_{0}\left(\frac{1}{n}\log\frac{\mu_{0}(x_{1}^{n})}{\mu_{1}(x_{1}^{n})}<\Delta\right)$
For $\Delta<D(\mu_{1}\|\mu_{0})$, let
$\lim_{n\to\infty}np_{n}(\Delta)=0\quad\text{and}\quad\sum_{n\geq 1}p_{n}(\Delta)<\infty$
For $\Delta<D(\mu_{0}\|\mu_{1})$, let
$\lim_{n\to\infty}nq_{n}(\Delta)=0\quad\text{and}\quad\sum_{n\geq 1}q_{n}(\Delta)<\infty$
II-A Previous Work
For the case where both pre-change $\mu_{0}$ and post-change $\mu_{1}$ distributions are known, Page proposed a test [4] which is given as follows.
Page’s CUSUM Test (Test 1) [4]: Starting with $n=1$, for samples $x_{1},x_{2},\ldots,x_{n}$, and $\gamma>1$, if
$\max_{k\leq n}\left(\log\mu_{1}(x_{k}^{n})-\log\mu_{0}(x_{k}^{n})\right)\geq\log\gamma,$ (2.18)
stop the test and decide that a change has occurred. Else, sample for $x_{n+1}$ and continue the test.
Jacob–Bansal Page CUSUM Test (Test 2)[24, 25]: A modification of Page’s CUSUM test was proposed for cases where the post-change distribution $\mu_{1}$ is unknown for i.i.d. sources and later extended for sources with memory satisfying (2.10),(2.16),(2.17). The modified CUSUM Test, where a strongly pointwise universal code is used to estimate the post-change distribution, is given as follows.
Starting with $n=1$, for samples $x_{1},x_{2},\ldots,x_{n}$, $\gamma>1$ and $\lambda<D(\mu_{1}\|\mu_{0})$, if
$\max_{k\leq n}\left(-L(x_{k}^{n})-\log\mu_{0}(x_{k}^{n})-n\lambda\right)\geq\log\gamma,$ (2.19)
stop the test and decide that a change has occurred. Else, sample for $x_{n+1}$ and continue the test.
Here, $\lambda$ was first used by the authors in order to ensure a negative drift in the test statistic in the pre-change regime, as is the case in the classical Page’s test in the parametric setup.
An auxiliary stopping time for this test is defined as
$N_{0}(\alpha)=\inf\left\{n:-L(x_{1}^{n})-\log\mu_{0}(x_{1}^{n})-n\lambda\geq-\log\alpha\right\}.$ (2.20)
An error occurs when Test 2 stops even though no change in distribution has occurred.The probability of error for the test is calculated using the stopping time $N_{0}(\alpha)$.
Malik–Jacob-Bansal Page CUSUM Test (Test 3) [26]: Since JB-Page’s CUSUM test requires the knowledge of pre-change distribution, a modification of this test was proposed for cases where the pre-change distribution $\mu_{0}$ is unknown for i.i.d. sources. The modified CUSUM Test, where a training sequence is used to estimate the pre-change distribution, is given as follows.
Test($\gamma>1$$\lambda>0$):Start with $n=1$,for samples $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$.If:
$\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}({x}_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma,$ (2.21)
Stop the test and decide that a change has occurred. Otherwise, sample for $x_{n_{0}+n+1}$ and continue the test.
Readers can refer to [26] to understand why universal code cannot be used to estimate pre-change distribution.
III Modified CUSUM Test For Markov
We consider stationary and ergodic mutually independent first order Markov sources $\mu_{0}$ and $\mu_{1}$ satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$ as well as codes satisfying (2.7) for analysis which can be generalised for any finite order Markov sources satisfying these conditions.We also assume that if $\mu_{0}\in\mathcal{M}_{0}$ and $\mu_{1}\in\mathcal{M}_{1}$ then $\min_{\begin{subarray}{c}\mu_{0}\in\mathcal{M}_{0}\\ \mu_{1}\in\mathcal{M}_{1}\end{subarray}}D(\mu_{1}\|\mu_{0})=\lambda_{0}>0$.
III-A Estimating Pre-change Distribution
For a stream of source symbols with no change up to $n_{0}$, the test has two stages. In the first stage, we estimate the prechange distribution using the first $n_{0}$ symbols, written as $x_{1},x_{2},\dots,x_{n_{0}}$ and the subsequent symbols as $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$. The prechange distribution estimate, $\hat{\mu}_{0}$, is calculated from initial sample of $n_{0}$ symbols as,
$\hat{\mu}_{0}(b,a)=\frac{N(b,a)\mid x_{1}^{n_{0}}}{n_{0}-1}$
$\hat{\mu}_{0}(b)=\frac{N(b)\mid x_{1}^{n_{0}}}{n_{0}}$
$\hat{\mu}_{0}(x_{1}^{n_{0}})=\frac{\hat{\mu}_{0}(x_{1},x_{2})\hat{\mu}_{0}(x_{2},x_{3})\dots\hat{\mu}_{0}(x_{n_{0}-1},x_{n_{0}})}{\hat{\mu}_{0}(x_{2})\dots\hat{\mu}_{0}(x_{n_{0}})}$
$\hat{\mu}_{0}(x_{1}^{n_{0}})=\hat{\mu}_{0}(x_{1})\prod_{i=2}^{n_{0}}\hat{\mu}_{0}(x_{i}/x_{i-1})$
Here, $N(b)\mid x_{1}^{n_{0}}$ represents the number of times the symbol $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $N(b,a)\mid x_{1}^{n_{0}}$ represents the number of times the symbols $b,a$ jointly appear in the sequence $x_{1}^{n_{0}}$.Also, $\hat{\mu}_{0}(b)$ is the fraction of times $b$ appears in the sequence $x_{1}^{n_{0}}$, whereas $\hat{\mu}_{0}(b,a)$ is the fraction of times $b,a$ appear in the sequence $x_{1}^{n_{0}}$ jointly.
We assume that $n_{0}$ is large enough so that each symbol $b\in\mathcal{X}$ which has non-zero probability under ${\mu}_{0}$ has appeared at least once in $x_{1}^{n_{0}}$ and symbol pair $(b,a)\in\mathcal{X}^{2}$ which have non-zero probability under ${\mu}_{0}$ have appeared at least once in $x_{1}^{n_{0}}$ (else $\hat{\mu}_{0}$ is not an accurate estimate of ${\mu}_{0}$).
Since the process is stationary,irreducible and aperiodic Markov Chain satisfying Berk’s Mixing Conditions,by Ergodic Theorem we can say that there exists an $n_{0}$ such that,
$\mu_{0}\left(\left\{x_{1}^{n_{0}}:\big|\hat{\mu}_{0}(a\mid b)-\mu_{0}(a\mid b)\big|>\delta\right\}\right)\leq\epsilon,$
for one or more $(b,a)\in\mathcal{X}^{2}.$
$\mu_{0}\left(\{x_{1}^{n_{0}}:|\hat{\mu}_{0}(b)-\mu_{0}(b)|>\delta^{\prime}\}\right)\leq\epsilon^{\prime},$
for one or more $b\in\mathcal{X}$
III-B Test
Using the estimate $\hat{\mu}_{0}$ and extending the Modified CUSUM Test for Markov sources by defining it as,
Test($\gamma>1$$\lambda>0$): Start with $n=1$ for samples $x_{n_{0}+1},x_{n_{0}+2},\dots,x_{n_{0}+n}$.If:
$\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}({x}_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma$
Stop the test and decide that a change has occurred. Otherwise, sample for $x_{n_{0}+n+1}$ and continue the test. The upward drift rate of this test is $D(\mu_{1}||\hat{\mu}_{0})-\lambda$ and the downward drift rate of this test is $\lambda-D(\mu_{0}||\hat{\mu}_{0})$. Hence, for this test to work, need:
$D(\mu_{0}||\hat{\mu}_{0})<\lambda<D(\mu_{1}||\hat{\mu}_{0})$
which implies,
$D(\mu_{0}||\hat{\mu}_{0})<D(\mu_{1}||\hat{\mu}_{0})$
Auxiliary stopping time for this modified test is defined as,
$N(\alpha)=\inf\left\{n:\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq-\log\alpha\right\}$ (3.10)
III-C Behaviour Under Prechange Distribution
Introducing the following notation to study the expected behaviour of the auxillary stopping time $N(\alpha)$ under prechange distribution $\mu_{0}$ and postchange distribution $\mu_{1}$ with respect to threshold parameter $\alpha$.
$f_{n}(x_{n_{0}+1}^{n_{0}+n})=\frac{\log\frac{\mu_{0}(x_{n_{0}+1}^{n_{0}+n})}{\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})}}{n}$
$l_{n}(x_{1}^{n_{0}+n})=-L(x_{n_{0}+1}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+1}^{n_{0}+n})-n\lambda$
$B_{n}\coloneq\{x_{1}^{n_{0}+n}:l_{k}(x_{1}^{n_{0}+k})<-\log\alpha,\;1\leq k<n,\;l_{n}(x_{1}^{n_{0}+n})\geq-\log\alpha\}$
$B=\bigcup_{n=1}^{\infty}B_{n}$
$E_{n}\coloneq\left\{x_{1}^{n_{0}+n}:\begin{array}[]{l}\big|\hat{\mu}_{0}(a\mid b)-\mu_{0}(a\mid b)\big|<\delta\quad\forall\,(a,b)\in\mathcal{X}^{2},\\[6.0pt] \big|\hat{\mu}_{0}(b)-\mu_{0}(b)\big|<\delta^{\prime}\quad\forall\,b\in\mathcal{X}\end{array}\right\},$
$\text{where }\hat{\mu}_{0}\text{ is based on samples }x_{1}^{n_{0}}.$
$E=\bigcup_{n=1}^{\infty}E_{n}$
Lemma 2: If $x_{1}^{n_{0}+n}\in E_{n}$, then $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ is bounded above for all $n$ as $f_{n}\leq\log\beta$ where $\beta=\beta^{\prime}\cdot\beta^{\prime\prime},\beta^{\prime}=\frac{\mu_{0}(a\mid b)}{\mu_{0}(a\mid b)-\delta},\beta"=\frac{\mu_{0}(b)}{\mu_{0}(b)-\delta^{\prime}}$ and $b,a$ is such that $b,a$ is the minimum probability symbol pair and $b$ is such that $b$ is the minimum probability symbol.As $\delta\to 0$$\delta^{\prime}\to 0$ and $n_{0}\to\infty$, we have $\beta\to 1$.
Proof:As $x_{i}$’s follow first order markov process satisfying (2.10),(2.16),(2.17), can write $f_{n}\big(x_{n_{0}+1}^{n_{0}+n}\big)$ as,
$f_{n}\big(x_{n_{0}+1}^{n_{0}+n}\big)=\sum_{(b,a)\in\mathcal{X}^{2}}\frac{N(b,a\mid x_{n_{0}+1}^{n_{0}+n})}{n}\cdot\log\frac{\mu_{0}(a\mid b)}{\widehat{\mu}_{0}(a\mid b)}$
$\sum_{b\in\mathcal{X}}\frac{N(b\mid x_{n_{0}+1})}{n}\cdot\log\frac{\mu_{0}(b)}{\widehat{\mu}_{0}(b)}$
As $x_{n_{0}+1}^{n_{0}+n}\in E_{n}$, for $x_{i}$’s sitting at any positions for $i=n_{0}+1,\ldots,n_{0}+n$,
$\frac{\mu_{0}(a\mid b)}{{\widehat{\ \mu}}_{0}(a\mid b)}<\frac{\mu_{0}(a\mid b)}{\mu_{0}(a\mid b)-\delta},\frac{\mu_{0}(b)}{{\widehat{\ \mu}}_{0}(b)}<\frac{\mu_{0}(b)}{\mu_{0}(b)-\delta^{\prime}}$
The terms on RHS are decreasing functions of $\mu_{0}(a\mid b)$ and $\mu_{0}(b)$ respectively, So if $b^{\prime}$ is the minimum probability symbol and $(b^{\prime},a^{\prime})$ is the minimum probability symbol pair then we can bound the above expression for $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ as (for $\delta\leq{\mu_{0}(a\mid b)}$ and $\delta^{\prime}\leq{\mu_{0}(b)}$):
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\log\left(\frac{\mu_{0}(a^{\prime}\mid b^{\prime})}{\mu_{0}(a^{\prime}\mid b^{\prime})-\delta}\right)+\log(\frac{\mu_{0}(b^{\prime})}{\mu_{0}(b^{\prime})-\delta^{\prime}})$
$f_{n}(x_{n_{0}+1}^{n_{0}+n})\leq\log(\beta^{\prime})+\log(\beta^{\prime\prime})=\log(\beta)$
Lemma 3:The probability of error (false alarm rate) for the proposed test is bounded above and given by
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{2^{\lambda-\log\beta}-1}+\epsilon_{0}$
Where $\epsilon_{0}\to 0$ and $\beta\to 1$ as $n_{0}\to\infty$.
Proof:Using the definition of $f_{n}(x_{n_{0}+1}^{n_{0}+n})$,
$\mu_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)=\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,nf_{n}(x_{n_{0}+1}^{n_{0}+n})}$
The probability of error is given as,
$\mu_{0}\!\left(N(\alpha)<\infty\right)=\mu_{0}(B)=\sum_{n=1}^{\infty}\mu_{0}\!\left(B_{n}\cap E_{n}\right)+\mu_{0}\!\left(B\cap(E)^{c}\right)$
$\sum_{n=1}^{\infty}\left(\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}\mu_{0}\!\left(x_{1}^{n_{0}+n}\right)\right)+\mu_{0}\!\left(B\cap(E)^{c}\right)$
As $B_{n}$ consists of only $n$-long sequences, we have $B_{n}\cap E=B_{n}\cap E_{n}$.For first term of (3.18) we have,
$\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}\mu_{0}\bigl(x_{1}^{n_{0}+n}\bigr)\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}\mu_{0}\bigl(x_{n_{0}+1}^{n_{0}+n}\bigr)$
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}(\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,nf_{n}(x_{n_{0}+1}^{n_{0}+n})}\bigr)$
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}(\hat{\mu}_{0}\!\left(x_{n_{0}+1}^{n_{0}+n}\right)\cdot 2^{\,n\log\beta}\bigr)$
$\leq\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}2^{\log\alpha-L(x_{n_{0}+1}^{n_{0}+n})\,-n(\lambda-\log\beta)}$
$\alpha\cdot 2^{\,-n(\lambda-\log\beta)}(\sum_{x_{1}^{n_{0}+n}\in B_{n}\cap E_{n}}2^{-L(x_{n_{0}+1}^{n_{0}+n})})$
$\leq\alpha\cdot 2^{\,-n(\lambda-\log\beta)}$
The second inequality comes from definition of $f_{n}(x_{n_{0}+1}^{n_{0}+n})$ while the third inequality is due to Lemma 2 as $x_{1}^{n_{0}+n}\in E_{n}$.The fourth inequality comes from definition of $B_{n}$ while the last inequality comes from Kraft’s inequality. As $P(A\cap B)\leq{P(B)}$,we have
$\mu_{0}(B\cap(E)^{c})\leq{\mu_{0}((E)^{c})}\leq{\epsilon_{0}}$
Combining both equations, we have
$\mu_{0}(N(\alpha)<\infty)\leq{\alpha}\sum_{n=1}^{\infty}2^{-n(\lambda-\log\beta)}+\epsilon_{0}$
$\mu_{0}(N(\alpha)<\infty)\leq\frac{\alpha}{2^{\lambda-\log\beta}-1}+\epsilon_{0}$
By comparing the probability of error expressions in Lemma 3 and for actual pre-change case (JB-Page CUSUM Test), we observe that the probability of error increases in the case of the empirical pre-change distribution test. The probability of error now becomes a function of $\log\beta$. For the probability of error to be bounded, we require:
$\lambda>\max(\log\beta,D(\mu_{0}||{\widehat{\mu}}_{0}))$
In contrast, for the actual pre-change distribution case, $\lambda>0$ would have sufficed. However, as $n_{0}\to\infty$, the lower bound $\max(\log\beta,D(\mu_{0}||{\widehat{\mu}}_{0}))$ and the error term $\epsilon_{0}$ can be made arbitrarily small.
Hence,
$\max(\log\beta,D(\mu_{0}||{\widehat{\mu}}_{0}))<\lambda<D(\mu_{1}||\hat{\mu}_{0})$
III-D Behaviour Under Postchange Distribution
Lemma 4: Let the conditions of Lemma 1 hold true.The proposed test will terminate with probability 1 under $\mu_{1}$ if $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$ i.e.,
$\mu_{1}(N(\alpha)<\infty)=1$
Proof:The proof of this lemma is the same as Lemma 5.4 of [25], replacing $\mu_{0}$ with ${\widehat{\mu}}_{0}$.The replacement is valid in this case as the behaviour of the test under $\mu_{1}$ is identical to that of a test where prechange distribution is $\mu_{0}$.
Theorem 2: Let $\mu_{0}$ and $\mu_{1}$ be stationary and ergodic finite order Markov sources satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$. For the Modified CUSUM Test (3.7) with codes satisfying (2.7) and stopping time $N(\alpha)$ (3.10) as $\alpha\to 0$, for $\lambda<D(\mu_{1}||{\widehat{\mu}}_{0})$, we have
${E_{1}}(N(\alpha)\,|\,x_{1}^{n_{0}})\sim\frac{|\log\alpha|}{D(\mu_{1}||{\widehat{\mu}}_{0})-\lambda}$
Proof:This theorem can be proved by arguing along the same lines as Theorem 5.7 of [25]
IV Performance Bounds
To study the performance of the proposed modified test, we use the following parameter to denote the stopping time of the test,
$M(\gamma)=\inf\left\{n:\max_{1\leq k\leq n}\left(-L(x_{n_{0}+k}^{n_{0}+n})-\log\hat{\mu}_{0}(x_{n_{0}+k}^{n_{0}+n})-n\lambda\right)\geq\log\gamma\right\}$
We work with conditional expectation while studying performance under $\mu_{1}$, conditioned on the initial $n_{0}$-length segment generated by $\mu_{0}$ as we used the empirical distribution $\hat{\mu}_{0}$ generated by this segment in place of $\mu_{0}$.
Consider the case where first m1m-1 samples are generated by $\mu_{0}$ and the rest of the samples come from $\mu_{1}$. In this case, the change point is $m$.Let $P_{m}$} and $E_{m}$} be the corresponding probability distribution and expectation for this case.Then,
$P_{m}\big(x_{n_{0}+1}^{\,n_{0}+n}\big)=\mu_{0}\big(x_{n_{0}+1}^{\,n_{0}+m}\big)\,\mu_{1}\big(x_{n_{0}+m}^{\,n_{0}+n}\big),$
due to the mutual independence of sources before and after the change [16, 25].
For a given stopping time N,Lorden introduced the following quantity (4.2) to serve as minimax criterion for quickness of reaction to a change [6].It states that,
$\overline{E_{1}}(S)=\sup_{m\geq 1}\,\text{ess sup}\,{E_{m}}\left[(S-m+1)^{+}\left|X_{1},\dots,X_{m-1}\right|\right]$
$\overline{E_{1}}(S)$ is the smallest bound on the average no.of differently distributed samples observed before stopping, guaranteed irrespective of behaviour of samples before the change. To study the behaviour of $\overline{E}_{1}(M(\gamma)\,|\,x_{1}^{n_{0}})$, first have to investigate the characteristics of ${E_{0}}(M(\gamma))$, which is the expected time after which the algorithm detects a change and stops the test for the no source change case. Ideally, this has to be $\infty$ as the algorithm should not detect a change when there isn’t any.
Theorem 3 [16]: Lorden’s Theorem was extended for stationary and ergodic processes $(\mu_{0},\mu_{1})$ satisfying Berk’s mixing conditions (2.16),(2.17) such that,
${\mu_{0}}(N<\infty)\leq\alpha$
And for $k=1,2,\dots$, let $N_{k}$ be the stopping variable obtained by applying $N$ to $X_{k},X_{k+1},\dots$, such that,
$N^{\ast}=\min\{N_{k}+k-1\mid k=1,2,\dots\}$
Then $N^{\ast}$ is an extended stopping variable with:
${E_{0}}(N^{\ast})\geq\frac{1}{2}\left(1+\frac{1}{\alpha}\right)>\frac{1}{2\alpha}$
And,
$\overline{E_{1}}(N^{\ast})\leq{E_{1}}(N)$
Define $S_{\gamma}$ to be the class of all sequential change detection tests’ stopping times $S$ satisfying (for $\gamma>0$):
${E_{0}}(S)\geq\frac{1+\gamma}{2}>\frac{\gamma}{2}$
Theorem 4 [16]: For two stationary ergodic and mutually independent stochastic processes with memory $\mu_{0},\mu_{1}$ satisfying (2.10) and Berk’s mixing conditions(2.16),(2.17), as $\gamma\to\infty$, we have,
$\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\sim\frac{\log\gamma}{D(\mu_{1}||\mu_{0})}$
We consider stationary ergodic and mutually independent finite order markov sources $\mu_{0},\mu_{1}$ satisfying (2.10),(2.16),(2.17)
[…]
${E_{0}}\left(M(\gamma)\right)\geq\frac{1}{2}+\frac{\gamma}{\frac{2}{2^{\lambda-\log\beta}-1}+2\epsilon_{0}\gamma}>\frac{\gamma}{\frac{2}{2^{\lambda-\log\beta}-1}+2\epsilon_{0}\gamma}$
and as $\gamma\to\infty$, we have,
$\overline{E_{1}}(M(\gamma)\mid x_{1}^{n_{0}})\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}$
Proof: Equation (4.9) follows from (3.15) of Lemma 3 and (4.5) of Theorem 3(restatement of Theorem 1 from [16]) Equation (4.10) follows from (4.6) of Theorem 3(restatement of Theorem 1 from [16]) and Theorem 2.
Theorem 6:Let $\mu_{0}$ and $\mu_{1}$ be stationary and ergodic finite order Markov sources satisfying (2.10),(2.16),(2.17) and $\mu_{1}\ll\mu_{0}$.For the Modified CUSUM Test (3.7), we can choose $\lambda$ and $\eta$ such that stopping time $M(\eta)=M\in S_{\gamma}$ for universal codes satisfying (2.7) and as $\gamma\to\infty$ we have,
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\inf_{s\in S_{\gamma}}\overline{E_{1}}(S)\quad\text{a.s.}\ [\mu_{0}]$
where $\vartheta$ can be chosen arbitrarily close to 0.
Proof:Let
$\eta=\gamma\left(\frac{1}{2^{\lambda-\log(\beta)}-1}+\epsilon_{0}\gamma\right).$
Now, if $\epsilon_{0}=\frac{1}{\gamma}$, then $\eta$ can be written as
$\eta=\gamma\left(\frac{1}{2^{\lambda-\log(\beta)}-1}+1\right).$
Consider $M=M(\eta)$, Then by Lemma 3 and Theorem 3 (or equivalently by Theorem 5), we have,
$E_{0}(M)>\frac{\gamma}{2}.$
[…]
${\lambda}=D(\mu_{1}\|\hat{\mu}_{0})-\frac{D(\mu_{1}\|\mu_{0})}{1+\vartheta}$
As $\gamma\to\infty$, we have $\epsilon_{0}\to 0$, hence $n_{0}\to\infty$, which yields
$D(\mu_{1}\|\hat{\mu}_{0})\to D(\mu_{1}\|\mu_{0})\quad\text{a.s.}\ [\mu_{0}].$
Also, as $n_{0}\to\infty$, we have
$\overline{E_{1}}(M\mid x_{1}^{n_{0}})\to\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\quad\text{a.s.}\ [\mu_{0}].$
From Theorem 5 we have,
$\overline{E_{1}}(M\mid x_{1}^{n_{0}})\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}$
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim\frac{\log\gamma}{D(\mu_{1}||\widehat{\mu_{0}})-\lambda}\quad\text{a.s.}\ [\mu_{0}]$
Substituting for $\lambda$,
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\frac{\log\gamma}{D(\mu_{1}||{\mu_{0}})}\quad\text{a.s.}\ [\mu_{0}]$
From Theorem 4 (restatement of Theorem 4 of [16]) we have,
$\inf_{S\in S_{\gamma}}\overline{E_{1}}(S)\sim\frac{\log\gamma}{D(\mu_{1}||\mu_{0})}$
Hence, as $n_{0}\to\infty$, we get
$\overline{E_{1}}(M\mid x_{1},x_{2},x_{3},...)\sim(1+\vartheta)\inf_{s\in S_{\gamma}}\overline{E_{1}}(S)\quad\text{a.s.}\ [\mu_{0}]$
V Conclusion and Future Work
In this paper, we have extended the results of 2022 for stationary and ergodic finite order Markov sources under suitable conditions.One can aim to generalise these results further under Lai’s framework [20] who has weakened the requirement of independence of observations before and after change to obtain an extension of Lorden’s result.The performance of proposed test can also be studied under other alternate criteria introduced by Lai [20].We have evaluated under Lorden’s criteria as our work extends [24, 25, 26] all of which are studied under Lorden’s criteria.A natural direction for future work would be to study under Pollack’s criteria[14] where variants for unknown distributions exist and compare performance.